If A (6,3) ,B(-3,5) ,C(4,-2) and D(x,3x) are four points If the ratio of area of triangle DBC and triangle ABC is 1:2 then value of x will be

Grade:12

3 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallvi

Find the area of triangle ABC

ΔABC =49/2


similerly find the area of triangle DBC

ΔDBC =7|2x-1|

now given ΔDBC/ΔABC   =1/2

  so   7|2x-1|/49/2  = 1/2

       2 |2x-1|/7 =1/2

         |2x-1| =7/4

      2x-1  =±7/4


+ve sign                                             -ve sign


2x-1 = 7/4                                        2x-1 =-7/4

x =11/8                                            x=-3/8


Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin

madhuri
6 years ago
in triangle PQR and triangleDEF ,angle P=angleE=30degrees,PQ/DE=PR/EF and angle F=75degrees show that PQ=PR
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to the question.
 
The area of triangle with vertices: (x1, y1), (x2, y2), (x3, y3) is
A = ½ | x1(y2 – y3) + x2(y3 – y1) + x3(y1 – y2) |
 
Using this we get the area of ΔABC = 49/2
 
similarly, we find the area of triangle DBC
ΔDBC =7|2x-1|
 
now given ΔDBC/ΔABC   = 1/2
7|2x – 1|/(49/2)  = 1/2
2 |2x – 1|/7 =1/2
|2x – 1| = 7/4
2x – 1 = ±7/4
 
Hence, x = 11/8 or –3/8
 
Thanks and regards,
Kushagra